2 + (A+3)/2 = 5.5 Need to solve for A. I know the answer, but need to know how it was done.
Very simple, just can't remember the proper way to do this from 10 years ago in High school. thanks

Answer by dkspoet(13)

2 + (A+3)/2 = 5.5
-2 + 2 + (A+3)/2 = 5.5 - 2
2*(A+3)/2 = 2*3.5
A+3-3 = 7-3
A = 4